# Pretraining with Random Noise for Fast and Robust Learning without Weight Transport

### v1

@@ B.3 Additional results with deeper network
Figure B.3: Training with random noise and data in a deeper network. (a) Three-layer network trained with random noise and data sequentially. (b) Test accuracy of the network under various training conditions. (c) Alignment angle of the weights $\mathbf{W}_{1}$ and synaptic feedback $\mathbf{B}_{1}$ in the first layer of the network. (d) Alignment angle of the weights $\mathbf{B}_{2}$ and synaptic feedback $\mathbf{W}_{2}$ in the second layer of the network.
B.4 Additional results with CIFAR-10
Table B.3: Parameters and settings used in the experiment.
Network architecture
Name Setting
Dimensions [3072, 100, 10]
Activation function ReLU
Random noise training
Name Setting
Number of examples $5\times 10^{5}$
Batch size 64
Learning rate 0.0001
Optimizer Adam
CIFAR10 data training
Name Setting
Number of training data $5\times 10^{3}$
Number of test data $5\times 10^{3}$
Epochs 100
Batch size 64
Learning rate 0.0001
Optimizer Adam
Figure B.4: Sample of the CIFAR-10 data [77] used in the subsequent data training process. The dataset contains images of ten classes, i.e. airplanes, cars, birds, cats, deer, dogs, frogs, horses, ships, and trucks.
Figure B.5: Effect of random noise training on subsequent CIFAR-10 data training. (a) Design of the CIFAR-10 classification task. (b) Test accuracy during CIFAR-10 data training. (c) Alignment angle of weights $\mathbf{W}_{1}$ and synaptic feedback $\mathbf{B}_{1}$ during random noise training and subsequent CIFAR-10 data training.
Figure B.6: Weight updates dynamic of random noise training and data training with CIFAR-10. (a) Trajectory of weights ($\mathbf{W}_{1}$) toward synaptic feedback matrix ($\mathbf{B}_{1}$) in the PCA (principal component analysis) latent space for random noise and CIFAR-10 data training. (b) Distance between weights ($\mathbf{W}_{1}$) and the synaptic feedback matrix ($\mathbf{B}_{1}$). (c) Order dependence of the trajectory of weights ($\mathbf{W}_{1}$). (d) Distance between weights ($\mathbf{W}_{1}$) and the synaptic feedback matrix ($\mathbf{B}_{1}$) for alternating orders of random and CIFAR-10 data training.
C.1 Network architecture and training details
Table C.1: Parameters and settings used in the experiment.
Network architecture
Name Setting
Dimensions [784, 100, 100, 10]
Activation function ReLU
Random noise training
Name Setting
Number of examples $5\times 10^{5}$
Batch size 64
Learning rate 0.0001
Optimizer Adam
C.2 Refinement of readout probability by random noise training
Figure C.1: Distribution of the readout probability. (a) Output probability of readout neurons measured in untrained and randomly trained networks. (b) Distribution of the readout probability of networks for various types of inputs: (1) Random noise, (2) MNIST data, (3) FMNIST data, and (4) KMNIST data.
C.3 Random noise training reduces the effective dimensionality of weights
Figure C.2: Effective dimensionality of weights in various layers. (a) The architecture of the network used in the experiment in which the weights of the three layers are analyzed. (b-d) Singular value spectrum of weights in each layer of the untrained and randomly trained networks. (e-g) Effective dimensionality of weights in each layer during random noise training.
C.4 Training neural networks with various training sizes
Table C.2: Parameters and settings used in the experiment.
Network architecture
Name Setting
Dimensions [784, 100, 100, 10]
Activation function ReLU
MNIST data training
Name Setting
Number of training data [100, 200, 400, 800, 1600]
Number of test data $1\times 10^{3}$
Epochs 500
Batch size 64
Learning rate 0.0001
Optimizer Adam
Figure C.3: Training process and generalization error for different training set sizes. (a) Training loss. (b) Test loss. (c) Training loss vs. test loss. Each column indicates a network trained with a different training set size, where gray and blue lines represent untrained and randomly pretrained networks, respectively.
Figure C.4: Training and test accuracy for different training set sizes. (a) Training accuracy. (b) Test accuracy. Each column represents the network trained with different training set sizes. Gray and blue lines represent the untrained and randomly pretrained networks, respectively.
C.5 Training neural networks with various network depths
Table C.3: Parameters and settings used in the experiment.
Network architecture
Name Setting
Input, hidden layer, output dimensions 784, 100, 10
Number of hidden layers [2, 3, 4, 5, 6]
Activation function ReLU
MNIST data training
Name Setting
Number of training data 1600
Number of test data $1\times 10^{3}$
Epochs 500
Batch size 64
Learning rate 0.0001
Optimizer Adam
Figure C.5: Training process and generalization error for different network depths. (a) Training loss. (b) Test loss. (c) Training loss vs. test loss. Each column represents a network trained with a different network depth, where gray and blue lines represent untrained and randomly pretrained networks, respectively.
Figure C.6: Training and test accuracy with different network depths. (a) Training accuracy. (b) Test accuracy. Each column represents a network trained with a different network depth, where gray and blue lines represent untrained and randomly pretrained networks, respectively.
C.6 generalization test for out-of-distribution
Figure C.7: Dataset used in the out-of-distribution generalization test with the transformed dataset. (a) MNIST dataset used in the training in-distribution case. (b-d) Transformed MNIST dataset used in the generalization test in the out-of-distribution case. (b) Translated MNIST dataset, where each image is randomly translated in the range of $[-5\%,5\%]$ of the input image size on the x- and y-axis. (c) Scaled MNIST dataset. Each image is randomly scaled in the range of $[0.8,1.2]$. (d) Rotated MNIST dataset. Each image is randomly rotated in the range of $[-25,25]$ degrees.
Figure C.8: Dataset used in the out-of-distribution generalization test with benchmark dataset. (a) MNIST dataset [52] used in the training the networks in the out-of-distribution case. (b) USPS dataset [78] resized from $(16\times 16)$ to $(28\times 28)$ used in the generalization test in the out-of-distribution case.
D.1 Network architecture and training details
Table D.1: Parameters and settings used in the experiment.
Network architecture
Name Setting
Dimensions [784, 100, 100, 10]
Activation function ReLU
Random noise training
Name Setting
Number of examples $5\times 10^{5}$
Batch size 64
Learning rate 0.0001
Optimizer Adam
Meta-loss measurement
Name Setting
K 10
Inner steps 10
Inner learning rate 0.001
Optimizer Adam
Various task adaption
Name Setting
Number of training data $5\times 10^{3}$
Number of test data $5\times 10^{3}$
Epochs 100
Batch size 64
Learning rate 0.0001
Optimizer Adam
D.2 Tasks used during the measurement of meta-loss
Figure D.1: Tasks used to measure the meta-loss. (a) Task 1: MNIST [52] classification task. (b) Task 2: Fashion-MNIST [55] classification task. (c) Task 3: Kuzushiji-MNIST [56] classification task.
Appendix E Experimental environment
E.1 Data availability
The datasets used in this study are publicly available. The MNIST dataset is available at http://yann.lecun.com/exdb/mnist/. The USPS dataset is available at https://www.csie.ntu.edu.tw/~cjlin/libsvmtools/datasets/multiclass.html#usps. The Fashion-MNIST dataset is available at https://github.com/zalandoresearch/fashion-mnist. The Kuzushiji-MNIST dataset is available at https://github.com/rois-codh/kmnist. The CIFAR-10 dataset is available at https://www.cs.toronto.edu/~kriz/cifar.html.
E.2 Software used
Python 3.11 (Python software foundation) with PyTorch 2.1 and NumPy 1.26.0 was used to perform the simulation and the analysis. SciPy 1.11.4 was used to perform the statistical test and analysis. The custom Python codes used in this work will be made available at the GitHub repository after the publication of the paper.
E.3 Computing resources
All simulations were performed on a computer with an Intel Core i7-11700K CPU and an NVIDIA GeForce RTX 1080 GPU. The simulation code was parallelized using PyTorch’s built-in parallelization to utilize the GPU resources efficiently.

### v2

@@ D.1 Network architecture and training details
Task distribution (Dataset) MNIST, F-MNIST, K-MNIST
Next, we compared untrained networks with networks pretrained using random noise in terms of their adaptation to each task. We copied the networks and adapted them to three tasks: MNIST, Fashion-MNIST, and Kuzushiji-MNIST. Detailed hyperparameters used in task adaptation can be found in Table S11.
D.2 Measurement of meta-loss
Figure S19: Tasks used to measure the meta-loss. (a) Task 1: MNIST [52] classification task. (b) Task 2: Fashion-MNIST [54] classification task. (c) Task 3: Kuzushiji-MNIST [59] classification task.
Figure S20: Diagram of meta-learning. Each colored arrow indicates the gradient update direction corresponding to a specific task. The gray arrow represents the overall learning direction of the meta-learning process. Each colored point denotes adaptation to a specific task from the meta-learned network.
We considered three tasks: MNIST classification, Fashion-MNIST classification, and Kuzushiji-MNIST classification (Figure S19) to measure the meta-loss and assess the fast adaptation of networks pretrained with random noise. The diagram of the meta-learning process (Figure S20) illustrates that optimizing the network to minimize the meta-loss enables quicker adaptation to task distributions. Notably, our results (Figure 6a) demonstrate that random noise pretraining efficiently reduces the meta-loss without the need for exposure to task-specific data. Additionally, the trajectory of weights in latent space during adaptation to various tasks (Figure 6b) follows a pattern similar to that depicted in the diagram. This highlights the role of random noise pretraining as a form of meta-learning that facilitates rapid adaptation.
Appendix E Our contributions
Our goal is to enhance current strategies in deep learning by drawing insights from brain function. There exists a significant accuracy gap between backpropagation and biologically plausible learning strategies that do not involve weight transport. While backpropagation is effective, it is computationally intensive, requiring dynamic memory access for weight transport (i.e., accessing forward weights in memory to compute backward updates). Pretraining with random noise and aligning forward-backward weights achieves similar outcomes without relying on weight transport. Our interest is not solely in attaining performance comparable to backpropagation, but rather in achieving this without weight transport. Feedback alignment (and random noise pretraining) does not require weight transport and relies exclusively on local information. In the context of energy-efficient neuromorphic chip engineering, feedback alignment is sometimes employed for learning, albeit with some performance trade-offs. Our results demonstrate that such sacrifices are minimal in our model. Overall, our findings provide insights into how to reach levels of learning efficiency similar to backpropagation while circumventing the need for weight transport.
Appendix F Experimental details
Software used. Python 3.11 (Python software foundation) with PyTorch 2.1 and NumPy 1.26.0 was used to perform the simulation and the analysis. SciPy 1.11.4 was used to perform the statistical test and analysis. The code used in this work is available at https://github.com/cogilab/Random.
Computing resources. All simulations were performed on a computer with an Intel Core i7-11700K CPU and an NVIDIA GeForce GTX 1080 GPU. The simulation code was parallelized using PyTorch’s built-in parallelization to utilize the GPU resources efficiently.